A convex mirror has a focal length of -27.3 cm. Find the magnification produced by the mirror when the object distance is (a) 9.1 cm and (b) 16.3 cm.

(a) m1 = ?

(b) m2 = ?
Answer #1

by mirror formula

1/v+1/u=1/f

(i) u=9.1 cm

f=-27.3 cm

we get v=-6.825

magnification=-v/u=0.75

(ii) u=16.3 cm

we get v=-10.21 cm

magnification=0.626